OpenStudy (anonymous):

As a train accelerates away from a station, it reaches a speed of 4.5m/s in 4.8s . If the train's acceleration remains constant, what is its speed after an additional 7.0s has elapsed?

OpenStudy (abhisar):

Hello @guiwuzhe456 ! Welcome to OpenStudy !

OpenStudy (abhisar):

Do you the equations of motion ?

OpenStudy (abhisar):

Using the equation v=u+at We cn frst find the acceleration of the train as 4.5=0+a*4.8 => a= 4.5/4.8=0.937 m/s^2 Now again using the same equation we can find the final final velocity as v=0+0.937*11.8 => v= ?

OpenStudy (anonymous):

v=11.0566